

********************************************************************
UARS WEEKLY STATUS REPORT - 22 Nov 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was shut down for the yaw-around on November 16 and will resume
solar scanning operations on November 22.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM web site has moved. It has a new URL. Access the PEM web site
using http://wwwpem.space.swri.edu. The PEM data from the GUVI conjunction
periods has moved to http://wwwpem.space.swri.edu/guvi-pem.html.
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 1993129 has been
reprocessed and is archived in the DAAC. HEPSA data for 1993074, 1993075,
and between 1993130 and 1993169 have been reprocessed and waiting DAAC archival.
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from the first conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes when both instruments saw aurora.
PEM has seen no aurora during the second conjunction period. Due to a
spacecraft anomaly, no PEM data was collected during the third conjunction
period.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 November 2002 (Orbit 61045, GMT Day 313) through
15 November 2002 (Orbit 61149, GMT Day 319).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 317/1740 61116 POWER ON
HALOE 317/1854 61117 POWER ON
SUSIM 317/2038 61118 POWER ON
SUSIM 319/2252 61149 POWER OFF
SOLSTICE 319/2252 61149 FULLTIME TO DAY-ONLY OPS
The beta angle ranged between the angles of 40.8 degrees to 16.0
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 19 November 2002 (DOY 323).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.28 to 1.60 degrees C. Additional
battery performance data for each battery on 08 November and 15
November 2002 are provided in Appendix D.
The clock error ranged from -33.7 to 02.0 msec during this report
period. The clock error of -33.7 was caused by a setting change made
on the TAC to reflect the correct method of time tagging the data.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 06
SOLSTICE = 02
HALOE = 01
SUSIM = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Four (4) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
On 13 November 2002, the FOT was re-initializing the UFLTR Processor
to reduce attitude biases. When sending a partial table load to set
the Mode 3 initialization flag, a non-nominal roll error was seen. It
turns out that the BASETAB used to set this flag also selects Earth
Sensor-3. In this mode the spacecraft uses pitch information from
Earth Sensor-2 and takes roll data from Earth Sensor-1, which was off
at the time. The last stale value in the Earth Senor-1 was 5
degree's. This caused the spacecraft to chase a 5-degree roll error
that didn't exist, which in turn caused a negative acquisition on the
subsequent pass. An omni pass was immediately scheduled, the proper
earth sensor selection was made (which corrected the attitude errors)
and all of the instruments were powered off due to insufficient power.
After which the UFLTR reinitialization took place successfully,
followed by instrument turn ons for SOLSTICE, HALOE, SUSIM and HRDI.
The reinitialization has shown lower residuals in the FSS and quieter
gyro biases, but there is still a minor roll error which is induced
during each orbit. The FOT is unsure of its cause as of yet and is in
the process of investigating this behavior. It may take several days
to determine what is causing this disturbance.
MEETINGS
None this report period.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 18 November 2002 (DOY 322).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 15 November 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
61088 315/1709 02-435 OBC BANK DUMPS
61113 317/1341 02-436 UFLTR UFLTR RE-INITIALIZATION
61116 317/1706 02-436 UFLTR UFLTR RE-INITIALIZATION
61116 317/1740 02-436 UFLTR UFLTR RE-INITIALIZATION - CLEANUP
61116 317/1740 02-437 SOLSTICE POWER ON
61117 317/1854 02-438 HALOE POWER ON
61117 317/1854 02-436 UFLTR UFLTR RE-INITIALIZATION - COMPLETED
61118 317/2038 02-439 SUSIM POWER ON
61118 317/2113 02-439 SUSIM POWER ON CONTIUNED
61119 317/2256 02-440 ACS GYRO BIAS RESET
61121 318/0220 02-440 ACS GYRO BIAS RESET
61125 318/0833 02-440 ACS GYRO BIAS RESET
61134 318/2218 02-440 ACS GYRO BIAS RESET
61145 319/1556 02-443 OBC/CLK CLK ADJ FR 0.0 TO -41.2 MSEC/DAY
61145 319/1556 02-444 MACS ESA-2 OFF
61148 319/2035 02-445 MACS GYRO BIAS NOISE UPDATE
61148 319/2110 02-440 MACS GYRO BIAS RESET
61149 319/2252 02-446 SUSIM POWER OFF
61149 319/2252/ 02-447 SOLSTICE FULLTIME TO DAY-ONLY OPS
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 15 November 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 15 November 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
60973 308/0605 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
61003 310/0602 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
61028 311/2131 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 15 November 2002
08 November 2002 - GMT Day 312 Beta = 44.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -151.2/-179.2 -3.05/-4.62 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-44.8 +5.47/+4.51 +27.2/-15.2 24.8 1.040 20.5
3 +22.4/-22.4 +4.51/+3.23 +26.4/-15.2 24.8 1.030 20.2
15 November 2002 - GMT Day 319 Beta = 16.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -156.8/-179.2 -3.36/-4.62 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-56.0 +5.79/+4.51 +24.8/-12.8 25.1 1.040 21.1
3 +22.4/-56.0 +4.83/+3.23 +23.6/-12.0 25.1 1.040 20.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
- None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 15 November 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-057 FOT 61108 317/0544 31741 GENERIC LATE ACQUISITION
02-058 FOT 61113 317/1341 N/A NEG ACQ / UFLTR ERROR
02-057 FOT 61113 317/1739 31741 GENERIC LATE ACQUISITION
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-053 FOT 60630 285/1008 30382 GENERIC LATE ACQUISITION
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-056 FOT 60775 295/0122 N/A OBC CRASH
02-057 FOT 61108 317/0544 31741 GENERIC LATE ACQUISITION
02-058 FOT 61113 317/1341 N/A NEG ACQ / UFLTR ERROR
02-057 FOT 61113 317/1739 31741 GENERIC LATE ACQUISITION